# Silent Speech Sentence Recognition with Six-Axis Accelerometers using Conformer and CTC Algorithm

### v1

@@ III-E Performance with respect to Different Channels
III-F Performance with respect to Different Axes
@@
The model was trained using six axes. Axes 1-3 are for acceleration data and axes 4-6 for gyroscope data. Single axis performance evaluation, shown in Fig. 5(d), reveals that gyroscope data generally outperformed accelerometer data, with accuracies of 92.39% to 93.39% versus 87.62% to 92.42%. Fig. 5(e) shows the accuracy with the use of more axes. It can be seen from Fig. 5(e) that a mean accuracy of 95.37% is achieved using two axes, while the accuracy is 95.57% when all six axes are used. This indicates that data from two axes may be sufficient for accuracy performance. Fig. 5(f) illustrates the accuracy performance using a pair of two axes. It is observed from Fig. 5(f) that the range of accuracy is from 94.04% to 97.07%.
IV Discussion
The presented work demonstrates a significant advancement in silent speech sentence recognition using six-axis accelerometers and a Conformer-based neural network with CTC algorithm. The proposed silent speech recognition system has the potential to significantly improve the quality of life for individuals with speech impairments, helping them regain the ability to communicate with others in normal lives. The portability of the six-axis accelerometer-based system makes it particularly suitable for real-world applications, ensuring that patients can rely on it in their daily lives. While the achieved accuracy of 97.17% is impressive, several aspects warrant further discussion and exploration.
The proposed method addresses the spatiotemporal correlation of facial motion acceleration data by utilizing convolutional layers to extract features from adjacent time and space domains. The Transformer is used for contextual learning, and the CTC algorithm for the interpretation and representation of word segmentation.
@@ IV Discussion
The strength of the proposed method lies in its ability to dynamically recognize arbitrary silent sentences by autonomous segmentation and identification based on a word database, without prior knowledge of the sentences. This overcomes the limitations of previous studies that treated word classification or treated sentences as a whole entirety similar to the word [22, 45]. The previous methods are only capable of static recognition and lack flexibility and expressiveness. The users of those approaches are limited to fixed databases and unable to convey extra meanings. This approach allows for a more versatile and practical SSI system, capable of recognizing a vast number of sentence combinations within the database. The patient can express whatever he/she wants by using different combinations of words to build a sentence, just like normal speaking.
The method demonstrates robust performance, excelling in recognizing easily confused words and adapting across individuals. Its few-shot learning capability allows quick adaptation to new users with minimal data. The compact, non-invasive six-axis accelerometers enhance portability and usability, making the system ideal for daily use by individuals with speech impairments.
@@
The few-shot learning shows promising results, although the performance is still affected by the habits of individual speakers. Further work is required to improve its generalization capability. Combining accelerometers with microphones could provide additional acoustic information to address this issue. Additionally, the relatively small dataset used in this study limits the model’s ability to handle complex linguistic phenomena. Expanding the database with diverse sentences and linguistic features could further enhance robustness and accuracy performance.
@@
V Conclusion
In this paper, a new method based on six-axis accelerometers has been presented for silent sentence recognition. The system is designed to let individuals with speech impairments regain the ability to communicate normally and enhance their expressive capabilities, focusing not only on word-level expression but also on complete sentence output. Six accelerometers are used to extract the motion features during the silent speaking period. A Conformer-based neural network integrated with the CTC algorithm was applied to translate mechanical motions into word and sentence transcriptions accurately. The average accuracy in sentence recognition and word classification is 97.17% based on the proposed method. The proposed method performs well on easily confused word recognition and shows robustness and generalization ability on the few-shot task.
This work is the first to recognize silent sentences dynamically and promotes the practical implementation and utilization of the system. The system is portable and applicable for patients in their daily usage due to the advantages of the accelerometers. In the future, the integration of accelerometers and other modalities, such as microphones, will be explored.
@@ Acknowledgments
The authors would like to thank Cancer Hospital Chinese Academy of Medical Sciences and volunteering participants for help in obtaining the dataset. We also thank Heng-Ming Tai and Kaigui Xie for their help on refining our manuscript.

### v2

@@ 3.5 The performance with respect to Different Channels
3.6 Performance with respect to Different Axes
The model was trained using six axes. Axes 1-3 are for acceleration data and axes 4-6 for gyroscope data. Single axis performance evaluation, shown in Fig. 5(d), reveals that gyroscope data generally outperformed accelerometer data, with accuracies of 92.39% to 93.39% versus 87.62% to 92.42%. This can be attributed to the biomechanical characteristics of facial and laryngeal articulation. During speech production, articulators such as jaw, undergo rotational motion patterns, pivoting at the temporomandibular joint.
@@ 3.6 Performance with respect to Different Axes
Fig. 5(e) shows the accuracy with the use of more axes. It can be seen from Fig. 5(e) that a mean accuracy of 95.37% is achieved using two axes, while the accuracy is 95.57% when all six axes are used. This indicates that data from two axes may be sufficient for accuracy performance. Fig. 5(f) illustrates the accuracy performance using a pair of two axes. It is observed from Fig. 5(f) that the range of accuracy is from 94.04% to 97.07%.
4 Discussion
@@
The strength of the proposed method lies in its ability to dynamically recognize arbitrary silent sentences by autonomous segmentation and identification based on a word database, without prior knowledge of the sentences. This overcomes the limitations of previous studies that treated word classification or treated sentences as a whole entirety similar to the word [22, 45]. The previous methods are only capable of static recognition and lack flexibility and expressiveness, while this approach allows for a more versatile and practical SSI system.
@@
Further work is required to improve its generalization capability. The relatively small dataset used in this study limits the model’s ability to handle complex linguistic phenomena. Expanding the database with diverse sentences and linguistic features could further enhance robustness and accuracy performance. Furthermore, instead of just the recognition of word sequences, the direct synthesis of speech can be examined. Based on the direct learning of speech by the accelerometer, we hope this system can enable generative speech communication in the future.
In addition to accuracy and model performance, the practical deployment of silent speech interfaces must consider sensor comfort. Future iterations could explore more ergonomic designs, such as thinner flexible substrates or skin-conformal electronics, to enhance wearability over extended periods.
Furthermore, continuous collection and processing of biometric motion data may raise privacy concerns. Ensuring data security, anonymization, and user control over data access will be essential for the responsible adoption of such systems in real-world healthcare or communication settings.
@@
5 Conclusion
In this paper, a new method has been presented for silent sentence recognition. The system focuses not only on word-level expression but also on complete sentence (sequence) output. Accelerometers combined with a Conformer-based neural network was applied to extract the facial motions and then translate mechanical motions into word and sentence transcriptions accurately. The average accuracy in sentence recognition and word classification is 97.17%. The proposed method performs well on easily confused word recognition and shows robustness and generalization ability on the few-shot task.
This work is the first to recognize silent sentences dynamically and promotes the practical implementation and utilization of the system. The system is portable and applicable for patients in their daily usage. In the future, the integration of accelerometers and other modalities, such as microphones, will be explored.
@@ Acknowledgments
The authors would like to thank Cancer Hospital Chinese Academy of Medical Sciences and volunteering participants for help in obtaining the dataset. We also thank Professor Heng-Ming Tai for his help on refining our manuscript.